OpenStudy (learner31):

Suppose p(a) = 0.76, p(b/a)= 0.30 and p(b/a bar)= 0.02. Find p(a and b), and p(a bar and b). use these to construct a probability table. Now use the table to find the following. a) p(b bar/ a) b) p( b bar/ a bar) c) p(b) d) p( b bar) e) p( a / b) f) p( a bar/ b) g) p( a/ b bar) h) p( a bar/ b bar)

OpenStudy (reemii):

Properties you can use \[ (1)\: P(\bar B | A) = 1 - P(B | A) \\ (2)\: P(A|B) = P(B|A)P(A)/P(B) (3)\: P(A|B) = P(A\cap B)/P(B) \] Not true though: \[ P(A|\bar B) = 1 - P(A|B)\]. This one would require you first to switch both events using (2): \[P(A|\bar B) = P(\bar B|A)P(A)/P(\bar B)\] and \(P(\bar B) = 1- P(B)\).

OpenStudy (reemii):

again, with better formatting. \[(1)\: P(\bar B | A) = 1 - P(B | A) \\ (2)\: P(A|B) = P(B|A)P(A)/P(B)\\ (3)\: P(A|B) = P(A\cap B)/P(B)\]

OpenStudy (reemii):

another one \[(4)\: P(A) = P(A \cap B) + P(A\cap \bar B)\]

OpenStudy (reemii):

with these, you can solve most or all of them

OpenStudy (learner31):

i am still not sure how to solve. can you help me?